# ACTISM: Threat-Informed Dynamic Security Modelling for Automotive Systems

### v2

@@ 4.1 Step 1: Identify assets
In this experience report, we reference the IVI HPC system DFD from Das et al. (2024) [4] to guide our identification of IVI system assets. This initial asset identification, listed below, serves as a foundational step for conducting further detailed analysis and security modelling. It is important to note that this DFD is an example and may not encompass newer types of attacks, such as dirty road attacks that exploit the fact that autonomous vehicles rely on sensors and cameras to interpret their environment. Attackers may place objects or patterns on the road that confuse these sensors, leading to incorrect navigation decisions or even accidents.
@@ 4.2 Step 2: Construct Data Flow Diagrams
In the next step, we construct and analyse DFDs to identify the vehicle trust boundary and potential threat and attack vectors. This process uses the Microsoft Threat Modelling Tool (Version 7.3.31026.3) along with the Automotive Threat Modelling Template [10] developed by the NCC Group. We conduct DFD analysis for both a typical automotive system [6] and specifically for the IVI system (Figure 2) to ensure thorough identification and analysis of potential threats and vulnerabilities.
@@
4.3 Step 3: Model threats
Next, we use the Microsoft Threat Modelling Tool to generate threats based on the previously constructed Data Flow Diagram (DFD) for the IVI system. The tool categorises these threats according to the STRIDE threat model automatically, providing an initial classification into categories such as Spoofing, Tampering, Repudiation, Information Disclosure, Denial of Service, and Elevation of Privilege.
@@ 4.3 Step 3: Model threats
We focus on the in-vehicle modules and data stores identified in the earlier step, ensuring that the generated threats are relevant to these components. After refining the list, we update threat descriptions to better align with the specific interfaces and nuances of the IVI system. A domain expert can perform this refinement relatively easily using the Microsoft Threat Modelling Tool, taking about 30 minutes for this example. The final curated list of threats for the IVI system is shown in Table 9 [8].
@@ 4.4 Step 4: Evaluate impact
    Each of the four parameters is assigned one of four impact levels: None, Low, Medium, or High. For the purposes of this experience report, we used the impact level definitions from [12]. Impact levels are typically assigned by automotive domain experts who have detailed knowledge of the automotive system, its vulnerabilities, and corresponding impacts on security objectives. In our running example, we assign impact levels based on our assessment of potential safety, operational, financial, and privacy consequences. This assessment is based on our understanding from the literature, and may differ from evaluations of experts in the field.
    Next, numerical values are assigned to the impact levels. These values are chosen in a logarithmic scale (0, 1, 10, 100) to effectively capture the exponential increase in the impact levels. This helps in clearly differentiating between the levels of impact related to the security objectives, and ensures that high-impact threats are given appropriate priority.
    Table 10 [8] categorises the impact levels and parameter values according to the security objectives.
@@
    $I_{\text{norm}}=\frac{w_{s}i_{s}+w_{o}i_{o}+w_{f}i_{f}+w_{p}i_{p}}{100*(w_{s}+w_{o}+w_{f}+w_{p})}$ (2)
@@
    $I=10i_{s}+i_{o}+i_{f}+i_{p}$ (3)
    $\small I_{\text{norm}}=\frac{10i_{s}+i_{o}+i_{f}+i_{p}}{1300}$ (4)
  • Assign impact rating: Each threat is assigned an impact rating based on the impact score $I_{\text{norm}}$ of each assessed threat, matched against thresholds tabulated in Table 2, as referenced from [12].
@@
    Impact Score Impact Rating
@@ 0.00 None
For illustration, we show how the impact rating of Threat ID 9146 from Table 9 [8] is derived using this methodology. In this specific threat, an adversary may reverse engineer the head unit firmware to find sensitive information, resulting in information disclosure from the IVI System’s local maps data store. Using the definitions of the impact assessment table (Table 10 [8]), we evaluate $i_{s}=0$, $i_{o}=0$, and $i_{f}=0$, as there are no injuries or discernible operational and financial effects resulting from the disclosure of information. However, disclosure of personally identifiable information or passwords from the data store can result in privacy violations leading to abuses such as impersonation of the victim to perform actions with stolen identities. As such, the impact level for privacy is assessed as Medium and $i_{p}=10$. Using Equation 2, $I_{\text{norm}}\ =0.0077$, which corresponds to a Low impact rating.
@@
4.5 Step 5: Model attacks
In this step, we use an attack tree to model attacker goals and map out attack paths associated with threat and attack vectors identified earlier using the Microsoft Threat Modelling Tool.
Since the vehicle driver is within the car trust boundary (Figure 2), we will assume the attacker personas and their goals as follows:
  • Persona A: A car thief interested in stealing sensitive data from the In-Vehicle Infotainment system.
@@ 4.5 Step 5: Model attacks
The attack tree depicted in Figure 3 illustrates the potential attack paths these two personas could employ to achieve their respective goals.
Figure 3: Attack tree for In-Vehicle Infotainment (IVI) system
@@ 4.6 Step 6: Evaluate attack feasibility
Next, we use CVSS version 3.1 to analyse and compute attack feasibility. CVSS provides a standardised method for assessing the severity of vulnerabilities based on their impact and exploitability factors, as described below.
@@ 5 Updating the Security Model After a Vulnerability Disclosure
ACTISM enhances security modelling by introducing a dynamic and iterative methodology, resulting in a threat-informed, evolving security model. This approach addresses evolving cybersecurity threats that necessitate the incorporation of new knowledge gained from security monitoring and threat intelligence throughout the system’s life cycle.
We now apply this methodology to a real-world example using the Tesla Electric Vehicle (EV) jail-break vulnerability disclosure [14]. Although this vulnerability is specific to the Tesla EV platform, ACTISM is designed to be applicable to all automotive systems (including non-EVs) and their relevant vulnerabilities.
After gathering technical details from the vulnerability disclosure, we review the IVI system security model developed in the earlier section and update the CVSS metrics as necessary.
For this vulnerability disclosure, there is no need to include additional assets in the security model or to revise the DFD (Figure 2) and attack tree (Figure 3). Although there is an opportunity to enhance the DFD and attack tree to include new attack vectors, such as cyber-physical attacks like keyless entry systems [15], we will not do so in this example. However, it is critical to reassess the CVSS metrics associated with Threat ID 9146, which correlates with the vulnerability disclosure.
@@
Besides attack feasibility, the impact assessment of Threat ID 9146 in the security model also needs updating. Using the definitions from the impact assessment table (Table 10 [8]), we estimate the impact on safety to increase from None to Low ($i_{s}=10$), as there is now a possibility of light and moderate injuries, should the attacker access and manipulate certain vehicle functions via the vulnerability. We also expect the operational impact to increase from None to Low ($i_{o}=10$), as there may be a degradation of primary vehicle functions. The financial impact is also expected to increase from None to Low ($i_{f}=10$), as the financial damage from injury claims may be substantial but not threatening the existence of the manufacturer. Lastly, the privacy impact is expected to increase from Low to High ($i_{p}=100$), as the privacy violation of multiple vehicle owners and drivers may lead to extensive media coverage and severe loss of market share. Using Equation 2 to compute the updated normalised impact score, $I_{\text{norm}}\ =0.1692$, which corresponds to a High impact rating (Table 8).
@@
6 Future directions
While ACTISM provides an approach to dynamically updating security models with both cyber and physical threats, it is not without its limitations. One of the primary challenges is how the accuracy of the model can be validated given the constantly evolving nature of threats and the potential for incomplete or inaccurate threat intelligence data. Moreover, while ACTISM’s adaptability to new threats is a strength, it can also pose a challenge in maintaining the models’ integrity and consistency over time. Despite these limitations, the potential future directions outlined below can address these challenges and enhance the robustness of ACTISM.
Ongoing threat intelligence gathering, including real-time monitoring and tracking of vulnerability disclosures, is crucial for keeping the security model current. Establishing a comprehensive repository of vulnerability attributes, stored in a standardised format such as XML or JSON, can enable automation and streamline the updating process. Additionally, formalising security models using Domain-Specific Languages (DSLs) like Automation Markup Language (AutomationML) and Meta Attack Language (MAL) can provide clearer descriptions of threats and vulnerabilities.
Incorporating adversary tactics from the MITRE ATT&CK knowledge base can further enhance these models by offering insights into real-world attack strategies. Moreover, leveraging machine learning approaches to prioritise vulnerabilities based on their impact, utilising historical data and threat intelligence, can significantly improve automotive cybersecurity.
These potential future directions can ensure that automotive cybersecurity remains resilient and adaptive to emerging cybersecurity challenges.
@@ 7 Conclusion
The cybersecurity landscape for automotive systems is rapidly evolving, with attackers increasingly targeting vulnerabilities that threaten data integrity, functionality, and safety. Traditional static security assessment methodologies, such as STRIDE, often fall short in addressing the dynamic and multi-faceted nature of these threats. There is a clear knowledge gap in the ability to model and respond to security threats in a dynamic, iterative manner tailored specifically for automotive systems.
@@
To illustrate the practical application of ACTISM, we developed a comprehensive security model for an IVI system. This process included identifying key assets, conducting thorough threat modelling, and performing risk assessments, ultimately resulting in a robust security framework. A real-world case study involving Tesla’s IVI system further demonstrated ACTISM’s ability to adapt to emerging threats and underscored its potential for broader applicability across various automotive technologies.
In conclusion, ACTISM addresses the critical knowledge gap in automotive cybersecurity by providing a proactive, adaptable approach to security modelling. This approach not only tackles current challenges but also prepares for the future landscape of automotive technology.

### v3

@@ 4.1 Step 1: Identify assets
In this experience report, we build on the IVI HPC system DFD from Das et al. (2024) [4] to guide our identification of IVI system assets. This initial asset identification, listed below, serves as a foundational step for conducting further detailed analysis and security modelling. While the referenced DFD provides a comprehensive framework for understanding IVI system components and interactions, we extend its scope to include emerging threats, such as dirty road attacks [14]. These attacks exploit the fact that autonomous and semi-autonomous vehicles rely heavily on sensors and cameras to interpret their environment. Attackers may place malicious objects, patterns, or signals on the road to confuse these sensors, leading to incorrect navigation decisions, system malfunctions, or even accidents. By incorporating such attack vectors, we aim to enhance the cyber-physical security modelling process and address the evolving threat landscape faced by modern automotives.
@@ 4.2 Step 2: Construct Data Flow Diagrams
In the next step, we construct and analyse DFDs to identify the vehicle trust boundary and potential threats and attack vectors, taking into consideration physical aspects. This process uses the Microsoft Threat Modelling Tool (Version 7.3.31026.3) along with the Automotive Threat Modelling Template [12] developed by the NCC Group. We conduct DFD analysis for both a typical automotive system [8] and specifically for the IVI system (Figure 2) to ensure thorough identification and analysis of potential threats and vulnerabilities.
@@
4.3 Step 3: Model threats (cyber & physical)
Next, we use the Microsoft Threat Modelling Tool to generate threats based on the previously constructed Data Flow Diagram (DFD) for the IVI system. The tool categorises these threats according to the STRIDE threat model automatically, providing an initial classification into categories such as Spoofing, Tampering, Repudiation, Information Disclosure, Denial of Service, and Elevation of Privilege. In addition to cyber threats, we also model physical threats, such as unauthorised physical access to the IVI system or tampering with hardware components.
@@ 4.3 Step 3: Model threats (cyber & physical)
We focus on the in-vehicle modules and data stores identified in the earlier step, ensuring that the generated threats are relevant to these components. After refining the list, we update threat descriptions to better align with the specific interfaces and nuances of the IVI system. A domain expert can perform this refinement relatively easily using the Microsoft Threat Modelling Tool, taking about 30 minutes for this example. The final curated list of threats for the IVI system is shown in Table 8 of our online appendix (see [10]).
@@ 4.4 Step 4: Evaluate impact
    Each of the four parameters is assigned an impact level: None, Low, Medium, or High, based on definitions from [11]. These levels are typically determined by automotive domain experts with in-depth knowledge of the system and its vulnerabilities. For this report, we assigned impact levels based on our assessment of potential consequences, informed by literature, though these may differ from expert evaluations.
    To quantify these levels, numerical values are assigned on a logarithmic scale (0, 1, 10, 100), reflecting the exponential increase in impact severity. This approach ensures clear differentiation between impact levels and prioritises high-impact threats effectively. Table 9 [10] categorises the impact levels and parameter values according to the security objectives.
@@
    $I_{\text{nsum}}=\frac{w_{s}i_{s}+w_{o}i_{o}+w_{f}i_{f}+w_{p}i_{p}}{100*(w_{s}+w_{o}+w_{f}+w_{p})}$ (2)
@@
    $I_{\text{sum}}=10i_{s}+i_{o}+i_{f}+i_{p}$ (3)
    $I_{\text{nsum}}=\frac{10i_{s}+i_{o}+i_{f}+i_{p}}{1300}$ (4)
  • Assign impact rating: Each threat is assigned an impact rating based on the impact score $I_{\text{nsum}}$, matched against thresholds tabulated in Table 1, as referenced from HEAVENS 2.0 [11]. The unequal ranges of the impact scores emphasise the importance of distinguishing between high-severity and critical-severity impacts, particularly in safety-critical systems like automotive systems. Similarly, the smaller ranges for lower-severity impacts reflect the fact that these scenarios are less critical and require less granularity in their assessment.
@@
    Impact Score ($I_{\text{nsum}}$) Impact Rating
@@ 0.00 None
For illustration, we show how the impact rating of Threat ID 9146 from Table 8 [10] is derived using this methodology. In this specific threat, an adversary may reverse engineer the head unit firmware to find sensitive information, resulting in information disclosure from the IVI System’s local maps data store. Using the definitions of the impact assessment table (Table 9 [10]), we evaluate $i_{s}=0$, $i_{o}=0$, and $i_{f}=0$, as there are no injuries or discernible operational and financial effects resulting from the disclosure of information. However, disclosure of personally identifiable information or passwords from the data store can result in privacy violations leading to abuses such as impersonation of the victim to perform actions with stolen identities. As such, the impact level for privacy is assessed as Medium and $i_{p}=10$. Using Equation 2, $I_{\text{nsum}}\ =0.0077$, which corresponds to a Low impact rating.
@@
4.5 Step 5: Model attacks (cyber & physical)
In this step, we use an attack tree to model attacker goals and map out attack paths associated with both physical and cyber threat and attack vectors. Since the vehicle driver is within the car trust boundary (Figure 2), we will assume the attacker personas and their goals as follows:
  • Persona A: A car thief interested in stealing sensitive data from the IVI system.
@@ 4.5 Step 5: Model attacks (cyber & physical)
Figure 3: Attack tree for an In-Vehicle Infotainment (IVI) system, illustrating the combination of physical and cyber tactics to achieve the attacker’s goals
The attack tree depicted in Figure 3 demonstrates how both personas could leverage a combination of physical and cyber tactics to achieve their objectives. One notable attack path included in the tree is the dirty road attack [14], where an attacker manipulates road signs (e.g. altering or obscuring traffic signals or directional signs) to mislead vehicle sensors. This physical manipulation could then be coupled with a cyber attack, such as exploiting a vulnerability in the IVI system to gain unauthorised access. By including both physical and cyber attack vectors, the attack tree underscores the importance of ACTISM’s holistic approach to security modelling, addressing not only cyber but also physical attack surfaces that could be exploited by malicious actors.
@@ 4.6 Step 6: Evaluate attack feasibility
Next, we use CVSS (Common Vulnerability Scoring System) version 3.1 to analyse and compute attack feasibility. CVSS provides a standardised method for assessing the severity of vulnerabilities based on their impact and exploitability factors, as described below. We have chosen to use CVSS 3.1 instead of the more recent CVSS 4.0 because CVSS 3.1 is currently more widely adopted across our tools, processes, and industry standards. This ensures consistency in scoring and compatibility with existing vulnerability management systems.
@@ 5 Updating the Security Model After a Vulnerability Disclosure
To demonstrate ACTISM’s dynamic and iterative methodology, we simulate a response to the Tesla Electric Vehicle (EV) jail-break vulnerability disclosure [15]. While this vulnerability is specific to Tesla EVs, ACTISM is designed to be applicable to all automotive systems, including non-EVs, and their associated vulnerabilities.
Using technical details from the disclosure, we review the IVI system security model developed earlier and update the CVSS metrics as needed. In this case, no additional assets, DFD updates (Figure 2), or attack tree revisions (Figure 3) are required. Although new attack vectors, such as cyber-physical attacks like dirty road attacks [14] or keyless entry systems [16], could be incorporated, we focus here on reassessing the CVSS metrics for Threat ID 9146, which aligns with the disclosed vulnerability. This process highlights ACTISM’s adaptability in responding to real-world threats.
@@
Besides attack feasibility, the impact assessment of Threat ID 9146 in the security model also needs updating. Using the definitions from the impact assessment table (Table 9 [10]), we estimate the impact on safety to increase from None to Low ($i_{s}=10$), as there is now a possibility of light and moderate injuries, should the attacker access and manipulate certain vehicle functions via the vulnerability. We also expect the operational impact to increase from None to Low ($i_{o}=10$), as there may be a degradation of primary vehicle functions. The financial impact is also expected to increase from None to Low ($i_{f}=10$), as the financial damage from injury claims may be substantial but not threatening the existence of the manufacturer. Lastly, the privacy impact is expected to increase from Low to High ($i_{p}=100$), as the privacy violation of multiple vehicle owners and drivers may lead to extensive media coverage and severe loss of market share. Using Equation 2 to compute the updated normalised impact score, $I_{\text{nsum}}\ =0.1692$, which corresponds to a High impact rating (Table 7).
@@
6 Practitioner Survey and Discussion
To assess the strengths of the ACTISM framework and identify areas for improvement, particularly in addressing emerging threats and enhancing usability in real-world automotive systems, feedback was gathered through a survey of 16 industry experts in Singapore. These experts included Chief Information Security Officers (CISOs) and cybersecurity practitioners specialising in security modelling and risk management. Representing a diverse range of industries across both private and public sectors, they provided a broad perspective on the evaluation of ACTISM. All respondents rated ACTISM as relevant, with a mean rating of 4.5 out of 5 based on a Likert scale, highlighting its potential as a robust framework for automotive cybersecurity. Furthermore, practitioners strongly agreed that ACTISM effectively integrates cyber and physical threat modelling (mean rating of 4.7 out of 5) and accurately identifies and models cyber-physical threats (mean rating of 4.1 out of 5). However, some respondents highlighted limitations in ACTISM, particularly its reliance on continuous threat intelligence updates to maintain effectiveness. They also noted challenges in implementing and using ACTISM compared to other security modelling frameworks. For automotive manufacturers with limited access to real-time cybersecurity threat intelligence feeds, as well as the absence of automation in the model workflow, these limitations could result in inconsistencies, false positives, and reduced operational efficiency over time.
To address these limitations and build on ACTISM’s strengths, we propose several future directions, many of which align with the recommendations provided by the experts. A key area for improvement is the need for adequate real-world or simulated testing to enhance the framework’s reliability. This would involve creating robust testing environments, including simulated attack scenarios, to evaluate ACTISM’s performance under realistic conditions. Such testing would also help validate the accuracy of the model, particularly given the constantly evolving nature of threats and the potential for incomplete or inaccurate threat intelligence data.
Another recommendation from the experts is the integration of ACTISM with complementary frameworks such as the Exploit Prediction Scoring System (EPSS). Integrating EPSS, which predicts the likelihood of a vulnerability being exploited, alongside the Common Vulnerability Scoring System (CVSS), which measures the severity of a vulnerability, could enhance ACTISM’s ability to prioritise and respond to threats effectively.
To further address the reliance on continuous threat intelligence updates, future work could focus on automating the model workflow. Establishing a comprehensive repository of vulnerability attributes, stored in a standardised format such as XML or JSON, would streamline the updating process and reduce the risk of inconsistencies. Additionally, formalising security models using Domain-Specific Languages (DSLs) like Automation Markup Language (AutomationML) and Meta Attack Language (MAL) could provide clearer descriptions of threats and vulnerabilities, improving the model’s usability and accuracy.
Incorporating adversary tactics from the MITRE ATT&CK knowledge base would further enhance ACTISM by offering insights into real-world attack strategies. This integration would enable the framework to better anticipate and mitigate sophisticated threats. Moreover, leveraging machine learning approaches to prioritise vulnerabilities based on their impact, utilising historical data and threat intelligence, could significantly improve the framework’s predictive capabilities and overall effectiveness.
These future directions will ensure that automotive cybersecurity remains resilient and capable of addressing emerging challenges in an increasingly complex threat landscape.
@@ 7 Conclusion
The cybersecurity landscape for automotive systems is rapidly evolving, with attackers increasingly targeting vulnerabilities that threaten data integrity, functionality, and safety. The siloed application of traditional security assessment methodologies using STRIDE and other frameworks often limits their ability to address the evolving and multi-faceted (cyber & physical) nature of these threats. There is a clear knowledge gap in the ability to model and respond to security threats in a dynamic, iterative manner tailored specifically for automotive systems.
@@
We applied ACTISM to Tesla’s In-Vehicle Infotainment system and showed that it can be practically used to model the dynamic nature of real-world cyber and physical threats in complex systems. Our experience report includes identifying key assets, conducting threat and attack modeling, and performing risk assessments for a complex system, in a dynamic, iterative manner. We also report the results of practitioners’ survey on the usefulness of ACTISM and its future directions. In conclusion, ACTISM bridges the critical gap in automotive cybersecurity by offering a proactive and adaptable approach to security modelling. This framework not only addresses current challenges but also prepares for the future, ensuring the resilience and safety of automotive systems in an ever-evolving threat landscape.